# Conformal Prediction with Corrupted Labels: Uncertain Imputation and Robust Re-weighting

### v2

@@ 1 Introduction
Modern machine learning models are increasingly deployed in high-stakes settings where reliable uncertainty quantification is essential. This need becomes even more critical when dealing with imperfect training data, which may be affected by noisy or missing labels. A common strategy to quantify prediction uncertainty is to construct prediction sets that cover the true outcome with a user-specified probability, e.g., 90%. Conformal prediction (CP[1] is a powerful framework for constructing theoretically valid prediction sets. Given a predictive model, CP utilizes labeled holdout calibration samples to compute prediction errors, which are then used to construct prediction sets for unseen test points. To provide a validity guarantee, this procedure requires that the training and test data are exchangeable, an assumption that does not hold in many real-world scenarios in which the observed data is corrupted.
@@
To illustrate the challenge of providing reliable inference under corrupted data, we conduct an experiment on the Medical Expenditure Panel Survey (MEPS) [2] dataset where the goal is to predict an individual’s medical utilization ($Y\in\mathcal{Y}$) given a set of features ($X\in\mathcal{X}$), such as race, income, medical conditions, and other demographic variables. We simulate a missing-at-random setup by randomly removing labels with a probability that depends on the features. Since we cannot use the samples with missing labels to compute the prediction error, we naively employ CP using only the observed data to construct prediction sets aiming to cover the true $Y$ with 90% probability. Figure 1 shows that this Naive CP fails to achieve the desired coverage due to the distributional shift induced by the missing labels.
@@
However, in this special case, the distributional shift is a covariate shift which we can account for using the method of weighted conformal prediction (WCP[3]. This method weights the data distribution by the labels’ likelihood ratio so that the train and test points will look exchangeable. Figure 1 reveals that WCP applied with the true weights attains the desired $1-\alpha=90\%$ coverage rate, as theoretically guaranteed in [3]. Nevertheless, WCP requires all test features $X$ to be observed to compute the weights. This assumption might not hold in practice, e.g., when an individual does not share sensitive attributes at test time, such as their income or race in our MEPS example, due to privacy concerns. In such cases, WCP is infeasible and cannot be employed as it is impossible to compute the weights.
Figure 1: Coverage rate obtained on the MEPS19 dataset.
@@
The goal of this work is to provide reliable inference under the setup of corrupted labels with missing features at test time. We refer to these features as privileged information (PI) [4]—additional information available during training but unavailable at test time. In our MEPS example, the privileged features are race, income, an individual’s rating of feeling, and more. In a classification task, the PI could include an annotator’s expertise level or the time taken to label an image. Although the PI may explain the annotation’s quality, it is unavailable at test time. Privileged conformal prediction (PCP[5] is a recent novel calibration scheme that builds on WCP to generate theoretically valid prediction sets without access to the test PI. However, PCP assumes access to the true weights of WCP, a requirement that might not hold in real-world applications. Indeed, Figure 1 shows that when applied with estimated weights, PCP does not achieve the nominal coverage level on the MEPS dataset. In this work, we focus precisely on this gap: what if the true weights are unavailable?
@@ 1.1 Our contribution
This work provides two key contributions. First, we analyze the robustness of PCP and WCP to inaccuracies in the approximated weights. We formally characterize the conditions under which these methods maintain valid coverage despite the errors of the weights. In contrast with prior work [6, 7, 8, 9] that deal with a worst-case analysis, our study reveals that PCP and WCP may construct prediction sets that attain the desired coverage rate even under significant errors in the weights, as demonstrated in empirical simulations in Section 3.1.2. Hence, the formulations developed in this work offer new theoretical guarantees and practical insights for these methods.
Second, we propose uncertain imputation (UI)—a novel calibration scheme that generates theoretically valid prediction sets in the presence of corrupted labels. In contrast to WCP and PCP, which assume that the weights can be estimated well from the PI, here, we assume that the clean labels can be estimated well from the PI [10, 11, 12]. Under several assumptions, we show how to impute corrupted labels in a way that preserves the uncertainty of the imputed labels. By leveraging recent results on label-noise robustness of conformal prediction [13, 14], we show that our uncertainty-preserving imputation guarantees the validity of our proposal. Indeed, Figure 1 demonstrates that our proposed UI achieves the desired coverage rate. Finally, we leverage the complementary validity conditions of PCP and UI, as well as CP, and propose combining all three into a triply robust calibration scheme (TriplyRobust) that constructs valid prediction sets as long as the assumptions of one of the methods are satisfied. Lastly, we conduct experiments on synthetic and real-world datasets to demonstrate the effectiveness of our proposed methods. Software implementing the proposed method and reproducing our experiments is available at https://github.com/Shai128/ui
@@ 1.2 Problem setup
Suppose we are given $n$ training samples $\{(X_{i},\tilde{Y}_{i},Z_{i},M_{i})\}_{i=1}^{n}$, where $X_{i}\in\mathcal{X}$ denotes the observed covariates, $\tilde{Y}_{i}\in\mathcal{Y}$ the observed, potentially corrupted, labels, $Z_{i}\in\mathcal{Z}$ the PI, and $M_{i}\in\{0,1\}$ is the corruption indicator. Specifically, if $M_{i}=0$ then $\tilde{Y}_{i}=Y_{i}$, where $Y_{i}$, is the clean, ground truth label. Otherwise, if $M_{i}=1$, then $\tilde{Y}_{i}$ is corrupted. For instance, in a missing response setup, if $M_{i}=1$ then $\tilde{Y}_{i}=\text{`}\texttt{NA}\text{'}$ ’. At test time, we are given the test features $X^{\text{test}}=X_{n+1}$, and our goal is to construct a prediction set $C(X^{\textup{test}})\subseteq\mathcal{Y}$ that covers the test response $Y^{\text{test}}=Y_{n+1}$ at a user-specified probability $1-\alpha$, e.g., 90%:
$\mathbb{P}(Y^{\textup{test}}\in C(X^{\textup{test}}))\geq 1-\alpha.$ (1)
This property is called marginal coverage, as the probability is taken over all samples $\{(X_{i},\tilde{Y}_{i},Y_{i},Z_{i},M_{i})\}_{i=1}^{n+1}$, which are assumed to be drawn exchangeably (e.g., i.i.d.) from $P_{X,\tilde{Y},Y,Z,M}$. The primary challenge in obtaining (1) is the distributional shift between the training data $\{(X_{i},\tilde{Y}_{i})\}_{i=1}^{n}$ and the test pair $(X_{n+1},Y_{n+1})$. In practice, naively applying CP on the corrupted data could lead to unreliable uncertainty estimates [15], or overly conservative prediction sets [13]. Moreover, calibrating using only the clean data introduces bias since the clean samples are drawn from $P_{X,Y\mid M=0}$, while the test distribution is $P_{X,Y}$. This bias might lead to undercoverage, as demonstrated in Figure 1.
To account for this bias, we assume the privileged information explains the corruption appearances, i.e., $(X,Y)\perp\!\!\!\perp M\mid Z$. We study the robustness of PCP [5] to inaccuracies of the weights under this assumption. We then consider an alternative setup where the clean label $Y$ can be estimated well from $Z$, and develop the method UI to construct prediction sets satisfying (1) even when the weights cannot be approximated accurately. Lastly, we combine all methods into a triply robust calibration scheme which enjoys the validity guarantees of both methods. See Table 1 in Appendix C.4 for a summary of our results.
@@ 2.1 Conformal prediction
Conformal prediction (CP[1] is a popular framework for generating prediction sets with a guaranteed coverage rate (1). It splits the dataset into a training set, denoted by $\mathcal{I}_{1}$, and a calibration set, denoted by $\mathcal{I}_{2}$. A learning model $\hat{f}$ is then trained on the training data and its performance is assessed on the calibration set using a non-conformity score function $\mathcal{S}(\cdot)\in\mathbb{R}$:
$S_{i}=\mathcal{S}(X_{i},Y_{i};\hat{f}),\forall i\in\mathcal{I}_{2}.$ (2)
For example, in regression problems, the score could be the absolute residual $\mathcal{S}(x,y;\hat{f})=|\hat{f}(x)-y|$, where $\hat{f}$ represents a mean estimator. The next step is computing the $(1+{1}/{|\mathcal{I}_{2}|})(1-\alpha)$-th empirical quantile of the calibration scores for the given $1-\alpha$ desired coverage level:
@@
This prediction set is guaranteed to achieve the desired marginal coverage rate, assuming that the calibration and test samples are exchangeable [1]. Since this is not the case in our setup, the next section introduces WCP, an extension of CP designed to handle covariate shifts.
@@ 2.2 Weighted conformal prediction
In the previous section, we noted that prediction sets constructed by CP might fail to achieve the desired coverage level due to the corruptions in the data. To overcome this issue, one could apply CP using only the scores of the uncorrupted samples, i.e., $\{\mathcal{S}(X_{i},\tilde{Y}_{i},\hat{f})\}_{i\in\mathcal{I}^{\text{uc}}_{2}}$, where $\mathcal{I}^{\text{uc}}_{2}=\{i\in\mathcal{I}_{2}:M_{i}=0\}$. Although these scores are computed using the true labels and are therefore accurate, considering only the uncorrupted samples induces a covariate shift between the calibration and test data. The weighted conformal prediction (WCP[3] method corrects this covariate shift by weighting the non-conformity scores using the likelihood ratio $w(z)=\frac{dP_{\text{test}}}{dP_{\text{train}}}(z)$, where $dP_{\text{test}}(z),dP_{\text{train}}(z)$ are the densities of the test and train probabilities, respectively. In this context, the weights can be expressed as $w(z)=\frac{\mathbb{P}(M=0)}{\mathbb{P}(M=0\mid Z=z)}$ [5]. Then, it extracts the quantile of the weighted distribution:
@@
The prediction set is constructed similarly to (4), except for using the threshold $Q^{\texttt{WCP}}(Z^{\text{test}})$. While this procedure is guaranteed to achieve the nominal coverage level, it cannot be applied directly since it relies on access to the test PI $Z^{\text{test}}$, which is unavailable in our framework.
@@ 2.3 Conformal prediction with noisy labels
The works of [13, 14, 16, 17, 18, 19] explore the setup where CP is employed with noisy labels. The main conclusion of [13] is that CP remains valid label noise when the noise is dispersive, i.e., increasing the variability of the observed labels. Specifically in regression tasks, their analysis reveals that CP remains valid when the noise is symmetric and additive. Their study also explores different noise models and provides empirical evidence for the robustness of CP. Building on these results, in Section 3.2 we impute corrupted labels with noisy versions of the true ones in a way that leads to a valid coverage rate. Additional related work is given in Appendix B.
@@ 3 Methods
In this section, we split our study into two cases based on the role of the PI: either as an indicator of the corruption indicator $M$ or as a proxy for the label $Y$. In the first case, for example, $Z$ may represent an annotator’s expertise level, such that a lower value may correspond to a higher likelihood of label noise. We present PCP, a method for constructing reliable uncertainty sets in this setup, and analyze its robustness to inaccuracies in the estimated corruption probabilities. In the second setting, we assume that the PI serves as a proxy for the label itself, for example, $Z$ may be a high-resolution image or detailed clinical reports available only during training. For this case, we develop UI–a novel imputation technique that leverages $Z$ to generate theoretically valid uncertainty sets.
@@ 3.1.1 Privileged conformal prediction
The PCP [5] procedure begins by partitioning the data into a training set, $\mathcal{I}_{1}$, and a calibration set, $\mathcal{I}_{2}$. Subsequently, a predictive model $\hat{f}$ is trained on the training set, and a non-conformity score is computed for each sample in the calibration set: $S_{i}=\mathcal{S}({X}_{i},\tilde{Y}_{i};\hat{f}),\forall i\in\mathcal{I}_{2}$. We also compute the likelihood ratio between the training and test distributions to calculate the weight for each sample $i$: $w_{i}:=\frac{\mathbb{P}(M=0)}{\mathbb{P}(M=0\mid Z=Z_{i})}$. Next, we treat each calibration point in $\mathcal{I}_{2}$ as a test point and apply WCP as a subroutine using the uncorrupted calibration samples to derive the score threshold $Q(Z_{i})$ for the $i$-th sample. The final test score threshold, denoted as $Q^{\texttt{PCP}}$, is then defined as the $(1-\beta)$-th empirical quantile of the calibration thresholds $\{Q(Z_{i})\}_{\{i\in\mathcal{I}_{2}\}}$:
@@
where $\beta\in(0,\alpha)$ is a pre-defined level, such as $\beta=0.05$. Finally, for a new test input $X^{\text{test}}$, the prediction set for $Y^{\text{test}}$ is constructed as follows: $C^{\texttt{PCP}}(X^{\text{test}})=\left\{y:\mathcal{S}(X^{\text{test}},y,\hat{f})\leq Q^{\texttt{PCP}}\right\}$.We now present [5, Theorem 1], stating the prediction sets constructed by PCP are valid.
Theorem 1 (Validity of PCP [5]).
Suppose that $\{({X}_{i},{Y}_{i},\tilde{Y}_{i},Z_{i},M_{i})\}_{i=1}^{n+1}$ are exchangeable, $Y\perp\!\!\!\perp M\mid Z$, and $P_{Z}$ is absolutely continuous with respect to $P_{Z\mid M=0}$. Then, the prediction set $C^{\texttt{PCP}}(X^{\textup{test}})$ constructed by PCP achieves the desired coverage rate: $\mathbb{P}(Y^{\textup{test}}\in C^{\texttt{PCP}}(X^{\textup{test}}))\geq 1-\alpha.$

### v3

# arXiv:2505.04733v3[cs.LG] 26 Feb 2026

@@ 1 Introduction
Modern machine learning models are increasingly deployed in high-stakes settings where reliable uncertainty quantification is essential. This need becomes even more critical when dealing with imperfect training data, which may be affected by noisy or missing labels. A common strategy to quantify prediction uncertainty is to construct prediction sets that cover the true outcome with a user-specified probability, e.g., 90%. Conformal prediction (CP(Vovk et al., 2005) is a powerful framework for constructing theoretically valid prediction sets. Given a predictive model, CP utilizes labeled holdout calibration samples to compute prediction errors, which are then used to construct prediction sets for unseen test points. To provide a validity guarantee, this procedure requires that the training and test data are exchangeable, an assumption that does not hold in many real-world scenarios in which the observed data is corrupted.
@@
To illustrate the challenge of providing reliable inference under corrupted data, we conduct an experiment on the Medical Expenditure Panel Survey (MEPS) (meps_, 19) dataset where the goal is to predict an individual’s medical utilization ($Y\in\mathcal{Y}$) given a set of features ($X\in\mathcal{X}$), such as race, income, medical conditions, and other demographic variables. We simulate a missing-at-random setup by randomly removing labels with a probability that depends on the features. Since we cannot use the samples with missing labels to compute the prediction error, we naively employ CP using only the observed data to construct prediction sets aiming to cover the true $Y$ with 90% probability. Figure 1 shows that this Naive CP fails to achieve the desired coverage due to the distributional shift induced by the missing labels.
@@
However, in this special case, the distributional shift is a covariate shift which we can account for using the method of weighted conformal prediction (WCP(Tibshirani et al., 2019). This method weights the data distribution by the labels’ likelihood ratio so that the train and test points will look exchangeable. Figure 1 reveals that WCP applied with the true weights attains the desired $1-\alpha=90\%$ coverage rate, as theoretically guaranteed in (Tibshirani et al., 2019). Nevertheless, WCP requires all test features $X$ to be observed to compute the weights. This assumption might not hold in practice, e.g., when an individual does not share sensitive attributes at test time, such as their income or race in our MEPS example, due to privacy concerns. In such cases, WCP is infeasible and cannot be employed as it is impossible to compute the weights.
Figure 1: Coverage rate obtained on the MEPS19 data.
@@
The goal of this work is to provide reliable inference under the setup of corrupted labels with missing features at test time (Collier et al., 2022; Wu et al., 2021; Ortiz-Jimenez et al., 2023). We refer to these features as privileged information (PI) (Vapnik & Vashist, 2009)—additional information available during training but unavailable at test time. In our MEPS example, the privileged features are race, income, an individual’s rating of feeling, and more; see Appendix B for more practical examples of PI. Privileged conformal prediction (PCP(Feldman & Romano, 2024) is a recent novel calibration scheme that builds on WCP to generate theoretically valid prediction sets without access to the test PI. However, PCP assumes access to the true weights of WCP, a requirement that might not hold in real-world applications. Indeed, Figure 1 shows that when applied with estimated weights, PCP does not achieve the nominal coverage level on the MEPS dataset. In this work, we focus precisely on this gap: what if the true weights are unavailable?
@@ 1.1 Our contribution
This work provides two key contributions. First, we analyze the robustness of PCP and WCP to inaccuracies in the approximated weights. We formally characterize the conditions under which these methods maintain valid coverage despite the errors of the weights. In contrast with prior work (Lei & Candès, 2021; Bhattacharyya & Barber, 2024; Gui et al., 2024; Marmarelis et al., 2024) that deal with a worst-case analysis, our study reveals that PCP and WCP may construct prediction sets that attain the desired coverage rate even under significant errors in the weights, as demonstrated in empirical simulations in Section 3.1.2. Hence, the formulations developed in this work offer new theoretical guarantees and practical insights for these methods.
Second, we propose uncertain imputation (UI)—a novel calibration scheme that generates theoretically valid prediction sets in the presence of corrupted labels. In contrast to WCP and PCP, which assume that the weights can be estimated well from the PI, here, we assume that the clean labels can be estimated well from the PI (Wu et al., 2021; Xu et al., 2021; Collier et al., 2022). Under several assumptions, we show how to impute corrupted labels in a way that preserves the uncertainty of the imputed labels. By leveraging recent results on label-noise robustness of conformal prediction (Einbinder et al., 2023; Sesia et al., 2024), we theoretically show that our uncertainty-preserving imputation guarantees the validity of our proposal even when the weights are unreliable. Importantly, PCP might fail to achieve the nominal coverage level in such cases. Indeed, Figure 1 demonstrates that our proposed UI achieves the desired coverage rate while PCP does not.
Finally, we leverage the complementary validity conditions of PCP, UI, and CP, and propose combining all three into a triply robust calibration scheme (TriplyRobust) that constructs valid prediction sets when the assumptions of one of the methods are satisfied. Lastly, we conduct experiments on synthetic and real datasets to demonstrate the effectiveness of our proposed methods. Software implementing the proposed method and reproducing our experiments is available at https://github.com/Shai128/ui
@@ 1.2 Problem setup
Suppose we are given $n$ training samples $\{(X_{i},\tilde{Y}_{i},Z_{i},M_{i})\}_{i=1}^{n}$, where $X_{i}\in\mathcal{X}$ denotes the observed covariates, $\tilde{Y}_{i}\in\mathcal{Y}$ the observed, potentially corrupted, labels, $Z_{i}\in\mathcal{Z}$ the PI, and $M_{i}\in\{0,1\}$ is the corruption indicator. Specifically, if $M_{i}=0$ then $\tilde{Y}_{i}=Y_{i}$, where $Y_{i}$, is the clean, ground truth label. Otherwise, if $M_{i}=1$, then $\tilde{Y}_{i}$ is corrupted. For instance, in a missing response setup, if $M_{i}=1$ then $\tilde{Y}_{i}=\text{`}\texttt{NA}\text{'}$. At test time, we are given the test features $X^{\text{test}}=X_{n+1}$, and our goal is to construct a prediction set $C(X^{\textup{test}})\subseteq\mathcal{Y}$ that covers the test response $Y^{\text{test}}=Y_{n+1}$ at a user-specified probability $1-\alpha$, e.g., 90%:
$\mathbb{P}(Y^{\text{test}}\in C(X^{\text{test}}))\geq 1-\alpha.$ (1)
This property is called marginal coverage, as the probability is taken over all samples $\{(X_{i},\tilde{Y}_{i},Y_{i},Z_{i},M_{i})\}_{i=1}^{n+1}$, which are assumed to be drawn exchangeably (e.g., i.i.d.) from $P_{X,\tilde{Y},Y,Z,M}$. The primary challenge in obtaining equation 1 is the distributional shift between the training data $\{(X_{i},\tilde{Y}_{i})\}_{i=1}^{n}$ and the test pair $(X_{n+1},Y_{n+1})$. In practice, naively applying CP on the corrupted data could lead to unreliable uncertainty estimates (Barber et al., 2023), or overly conservative prediction sets (Einbinder et al., 2023). Moreover, calibrating using only the clean data introduces bias since the clean samples are drawn from $P_{X,Y\mid M=0}$, while the test distribution is $P_{X,Y}$. This bias might lead to undercoverage, as demonstrated in Figure 1.
To account for this bias, we assume the privileged information explains the corruption appearances, i.e., $(X,Y)\perp\!\!\!\perp M\mid Z$. We study the robustness of PCP (Feldman & Romano, 2024) to inaccuracies of the weights under this assumption. We then consider an alternative setup where the clean label $Y$ can be estimated well from $Z$, and develop the method UI to construct prediction sets satisfying equation 1 even when the weights cannot be approximated accurately. Lastly, we combine all methods into a triply robust calibration scheme which enjoys the validity guarantees of both methods. See Table 1 in Appendix C.4 for a summary of our results.
@@ 2.1 Conformal prediction
Conformal prediction (CP(Vovk et al., 2005) is a popular framework for generating prediction sets with a guaranteed coverage rate equation 1. It splits the dataset into a training set, denoted by $\mathcal{I}_{1}$, and a calibration set, denoted by $\mathcal{I}_{2}$. A learning model $\hat{f}$ is then trained on the training data and its performance is assessed on the calibration set using a non-conformity score function $\mathcal{S}(\cdot)\in\mathbb{R}$: $S_{i}=\mathcal{S}(X_{i},Y_{i};\hat{f}),\forall i\in\mathcal{I}_{2}$. For example, in regression problems, the score could be the absolute residual $\mathcal{S}(x,y;\hat{f})=|\hat{f}(x)-y|$, where $\hat{f}$ represents a mean estimator. The next step is computing the $(1+{1}/{|\mathcal{I}_{2}|})(1-\alpha)$-th empirical quantile of the calibration scores for the nominal coverage level:
@@
This prediction set is guaranteed to achieve the desired marginal coverage rate, assuming that the calibration and test samples are exchangeable (Vovk et al., 2005). Since this is not the case in our setup, the next section introduces WCP, an extension of CP designed to handle covariate shifts.
@@ 2.2 Weighted conformal prediction
In the previous section, we noted that prediction sets constructed by CP might fail to achieve the desired coverage level due to the corruptions in the data. To overcome this issue, one could apply CP using only the scores of the uncorrupted samples, i.e., $\{\mathcal{S}(X_{i},\tilde{Y}_{i},\hat{f})\}_{i\in\mathcal{I}^{\text{uc}}_{2}}$, where $\mathcal{I}^{\text{uc}}_{2}=\{i\in\mathcal{I}_{2}:M_{i}=0\}$ are the indexes of the uncorrupted calibration samples. Although these scores are computed using the true labels and are therefore accurate, considering only the uncorrupted samples induces a covariate shift between the calibration and test data. The weighted conformal prediction (WCP(Tibshirani et al., 2019) method corrects this covariate shift by weighting the non-conformity scores using the likelihood ratio $w(z)=\frac{dP_{\text{test}}}{dP_{\text{train}}}(z)$, where $dP_{\text{test}}(z),dP_{\text{train}}(z)$ are the densities of the test and train probabilities, respectively. In this context, the weights can be expressed as $w(z)=\frac{\mathbb{P}(M=0)}{\mathbb{P}(M=0\mid Z=z)}$ (Feldman & Romano, 2024). Then, it extracts the quantile of the weighted distribution:
@@
The prediction set is constructed similarly to equation 3, except for using the threshold $Q^{\texttt{WCP}}(Z^{\text{test}})$. While this procedure is guaranteed to achieve the nominal coverage level, it cannot be applied directly since it relies on access to the test PI $Z^{\text{test}}$, which is unavailable in our framework.
@@ 2.3 Conformal prediction with noisy labels
The works of Einbinder et al. (2023); Sesia et al. (2024); Penso & Goldberger (2024); H. Zargarbashi et al. (2024); Penso et al. (2025); Bashari et al. (2025) explore the setup where CP is employed with noisy labels. The main conclusion of Einbinder et al. (2023) is that CP remains valid label noise when the noise is dispersive, i.e., increasing the variability of the observed labels. Specifically in regression tasks, their analysis reveals that CP remains valid when the noise is symmetric and additive. Their study also explores different noise models and provides empirical evidence for the robustness of CP. Building on these results, in Section 3.2, we impute corrupted labels with noisy versions of the true ones in a way that leads to a valid coverage rate. Additional related work is given in Appendix B.
@@ 3 Methods
In this section, we split our study into two cases based on the role of the PI: either as an indicator of the corruption variable $M$ or as a proxy for the label $Y$. In the first case, for example, $Z$ may represent an annotator’s expertise level, such that a lower value may correspond to a higher likelihood of label noise. We present PCP, a method for constructing reliable uncertainty sets in this setup, and analyze its robustness to inaccuracies in the estimated corruption probabilities. In the second setting, we assume that the PI serves as a proxy for the label itself, for example, $Z$ may be a high-resolution image or detailed clinical reports available only during training. For this case, we develop UI—a novel imputation technique that leverages $Z$ to generate theoretically valid uncertainty sets.
@@ 3.1.1 Privileged conformal prediction
The PCP (Feldman & Romano, 2024) procedure begins by partitioning the data into a training set, $\mathcal{I}_{1}$, and a calibration set, $\mathcal{I}_{2}$. Subsequently, a predictive model $\hat{f}$ is trained on the training set, and a non-conformity score is computed for each sample in the calibration set: $S_{i}=\mathcal{S}({X}_{i},\tilde{Y}_{i};\hat{f}),\forall i\in\mathcal{I}_{2}$. We also compute the likelihood ratio between the calibration and test distributions to calculate the weight for each sample $i$: $w_{i}:=\frac{\mathbb{P}(M=0)}{\mathbb{P}(M=0\mid Z=Z_{i})}$. Next, we treat each calibration point in $\mathcal{I}_{2}$ as a test point and apply WCP as a subroutine using the uncorrupted calibration samples to derive the score threshold $Q(Z_{i})$ for the $i$-th sample. The final test score threshold, denoted as $Q^{\texttt{PCP}}$, is then defined as the $(1-\beta)$-th empirical quantile of the calibration thresholds $\{Q(Z_{i})\}_{\{i\in\mathcal{I}_{2}\}}$:
@@
where $\beta\in(0,\alpha)$ is a pre-defined level, e.g., $\beta=0.05$. Finally, for a new test input $X^{\text{test}}$, the prediction set for $Y^{\text{test}}$ is constructed as follows: $C^{\texttt{PCP}}(X^{\text{test}})=\left\{y:\mathcal{S}(X^{\text{test}},y,\hat{f})\leq Q^{\texttt{PCP}}\right\}$.This prediction set is guaranteed to obtain a valid coverage rate, as stated next.
Theorem 1 (Validity of PCP (Feldman & Romano, 2024)).
Suppose that $\{({X}_{i},{Y}_{i},\tilde{Y}_{i},Z_{i},M_{i})\}_{i=1}^{n+1}$ are exchangeable, $Y\perp\!\!\!\perp M\mid Z$, and $P_{Z}$ is absolutely continuous with respect to $P_{Z\mid M=0}$. Then, the prediction set $C^{\texttt{PCP}}(X^{\textup{test}})$ achieves the desired coverage rate: $\mathbb{P}(Y^{\textup{test}}\in C^{\texttt{PCP}}(X^{\textup{test}}))\geq 1-\alpha.$